OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (texaschic101):

y - 4 = 3/4(x + 8) -- distribute through the parenthesis y - 4 = 3/4x + 6 -- add 4 to both sides y = 3/4x + 6 + 4 y = 3/4x + 10 In standard form we do not want fractions...so we need to multiply the entire equation by the LCD, which is 4. 4y = 3x + 40 -- subtract 3x from both sides -3x + 4y = 40 -- we do not want x to be negative, so we will multiply the equation by -1 3x - 4y = -40 <== standard form